### v1

# Beyond the Labels: Unveiling Text-Dependency in Paralinguistic Speech Recognition Datasets

Abstract
Paralinguistic traits like cognitive load and emotion are increasingly recognized as pivotal areas in speech recognition research, often examined through specialized datasets like CLSE and IEMOCAP. However, the integrity of these datasets is seldom scrutinized for text-dependency. This paper critically evaluates the prevalent assumption that machine learning models trained on such datasets genuinely learn to identify paralinguistic traits, rather than merely capturing lexical features. By examining the lexical overlap in these datasets and testing the performance of machine learning models, we expose significant text-dependency in trait-labeling. Our results suggest that some machine learning models, especially large pre-trained models like HuBERT, might inadvertently focus on lexical characteristics rather than the intended paralinguistic features. The study serves as a call to action for the research community to reevaluate the reliability of existing datasets and methodologies, ensuring that machine learning models genuinely learn what they are designed to recognize.
Index Terms:
Paralinguistic Traits, Speech Recognition, Cognitive Load, Emotion Recognition, Lexical Overlap, Machine Learning, Datasets, Text-Dependency
I Introduction
While the primary focus of speech recognition research gravitates towards Automatic Speech Recognition (ASR), the study of paralinguistic traits, such as cognitive load, physiological stress and emotions, remains a significant field too. These traits are of interest for applications, ranging from human-computer interaction to psychological research, and rely heavily on dedicated datasets. One such dataset is Cognitive Load with Speech and EGG (CLSE[1], commonly used for cognitive load recognition. Another key dataset is the  Interactive Emotional Dyadic Motion Capture (IEMOCAP[2], primarily employed for emotion recognition.
A prevailing assumption is that machine learning algorithms trained on these datasets learn to recognize paralinguistic traits based on observable physiological or psychological changes in speech production.
However, this paper challenges this assumption by revealing a critical oversight: we provide evidence of significant lexical correlation between the labels (e.g., cognitive load or emotion) and the uttered sentences within these datasets.
We extensively analyse CLSE and review the design of IEMOCAP to substantiate this claim. Given the implications of our findings, the paper serves as a call to action for the research community. It urges a reevaluation of existing datasets and methodologies to ascertain that machine learning systems are learning to recognize what they are designed to recognize.
II Background
II-A Datasets in Focus
Two key datasets facilitate paralinguistic studies:
  • CLSE (Cognitive Load with Speech and EGG): provides an approach to recognizing cognitive load by incorporating both speech recordings and Electroglottograph (EGG) signals.
  • IEMOCAP (Interactive Emotional Dyadic Motion Capture) [^1] https://sail.usc.edu/iemocap/index.html: covers about twelve hours of scripted and spontaneous dialogues. It captures detailed facial and hand movements to study a range of emotions.
II-B Machine Learning Approaches
Various machine learning algorithms, from classical UBM-iVector [3], through LSTM [4] to recent large pre-trained models like wav2vec [5], have been employed on these datasets. The general presumption is that these models are capturing paralinguistic features rooted in physiological or psychological changes [6, 7].
II-C The What and the How
It is crucial to note that in paralinguistic research, both the content (what) and manner (how) of speech carry significance. While the physiological and prosodic features often capture the how, the lexical choices, semantics, and syntax reveal insights into the what. Ignoring either aspect could result in an incomplete understanding of paralinguistic traits. Therefore, this study highlights the importance of acknowledging the lexical correlates when analyzing paralinguistic datasets.
III CLSE: Methodology
III-A Dataset and Task Description
Our primary study focuses on the main part of the CLSE database, termed as CLSE-Span, which employs the  Reading Span task (RSPAN[8] task. In RSPAN, participants are asked to validate sentences for their logical coherence while memorizing a sequence of letters. Sentences are randomly chosen from a closed set. The task is divided into sets, where each set contains varying trials of sentence validation and letter memorization. The database consists of 21 such sets with each participant producing 75 utterances, totaling 1800 utterances. The average utterance duration is 4 seconds. Labels for the cognitive load values are: low, medium and high load.
The metric for this study is the Unweighted Average Recall (UAR). UAR offers a balanced assessment of performance by averaging recall across all classes, making it particularly valuable for evaluating imbalanced datasets.
III-B Experimental setup
To explore the dataset, we initially employed ASR using Whisper [9] with a pre-trained multilingual large-v3 model to transcribe the utterances. Subsequently, we utilized Sentence Transformer embeddings as features. We used the Language-Agnostic BERT Sentence Embeddings (LaBSE) model [10].
LaBSE utilizes a dual-encoder architecture trained in two steps: first on 17 billion monolingual sentences via Masked Language Modeling, and then further refined on 6 billion translation pairs covering 109 languages. The final model is publicly accessible [^2] https://huggingface.co/sentence-transformers/LaBSE and was used to extract 768-dimensional sentence embeddings.
We then applied Agglomerative Hierarchical Clustering (AHC) on the generated utterance embeddings. After manual correction of minor clustering errors, we have identified 81 different sentences being spoken. Ground truth sentences are not available in the distributed CLSE version. The cluster size ranged between 12 and 26 utterances with majority size of 26 utterances. Given the cluster membership, we have assigned sentence ID to each utterance.
For each cluster, we identified the most frequent cognitive load label in the training set. These identified labels were then used as a proxy to predict the cognitive load of utterances in the test and validation sets based on their cluster identity, effectively bypassing the use of any machine learning model for classification. This approach revealed how well the dataset’s original labels correlated with the cluster-based predictions.
To further corroborate our hypothesis, we reshuffled the dataset splits based on cluster IDs. This process introduced speaker overlap but did not affect the cognitive load estimation, as each speaker experiences a full range of cognitive loads in the dataset.
IV CLSE:Results
IV-A Cluster-Based Classification: Impact of Lexical Overlap
We assessed the influence of lexical overlap on classification performance using both original and fixed (shuffled) CLSE data splits. The original CLSE data displayed good performance metrics, unfortunately revealing pattern in the validation set that confirmed lexical factors’ role.
TABLE I: Performance in Original vs Fixed CLSE Splits for Cluster-Based Classification
Original CLSE Fixed CLSE
Train 0.67 0.33
Validation 0.79 0.33
Test 0.57 0.33
After minimizing lexical redundancy in shuffled splits, the performance metrics notably decreased down to the chance level, as can be seen in Table I, further substantiating the lexical overlap’s impact on classification.
IV-B UBM-ivector system: ComPaRE 2014 replication
We have reached out to the winners of the ComPaRE 2014 [11] challenge and they helpfully provided us with the system setup of their winning system. Using their code, we were able to replicate and further improve on their results.
The UBM-iVector system uses a 64-component Universal Background Model (UBM) with 50 dimensional iVector extractor and Support Vector Machine (SVM) for classification. The optimal size for the system components were found by running grid search for UBM, iVector Extractor and SVM parameters. The details of the original system architecture can be found in [3].
Using the same approach as in Section IV-A, we have obtained results which shows significant degradation of the performance on validation and test set, see Table II.
TABLE II: Performance in Original vs Fixed CLSE Splits for ComPaRE2014 System
Split Original CLSE Fixed CLSE
Train 0.94 0.98
Validation 0.75 0.51
Test 0.64 0.51
IV-C HuBERT based system
To corroborate our findings with a more recent speech-based system, we used HuBERT [12] with pre-trained large-ll60k weights trained on 60k hours from the Libri-Light dataset [13] as a feature extractor. As a classifier, we used Attentive correlation pooling layer from [14]. The results of our experiments are shown in Table III.
TABLE III: Performance in Original vs Fixed CLSE Splits for HuBERT based System
Split Original CLSE Fixed CLSE
Train 0.82 0.76
Validation 0.76 0.64
Test 0.74 0.52
The results show the same trend as seen in the text-based and UBM-iVector based systems. Accros the board, removing the lexical context hurts the performance with at least absolute 12% UAR degradation.
V IEMOCAP: Methodology
V-A Dataset and Task Description
As a secondary experiment, we have analyzed IEMOCAP dataset. It comprises 151 dialogue videos, featuring a pair of speakers in each session, resulting in a total of 302 individual videos with aproximately 12 hours of data. Ten actors were recorded in dyadic sessions (5 sessions with 2 subjects each). They were asked to perform three selected scripts with clear emotional content. In addition to the scripts, the subjects were also asked to improvise dialogs in hypothetical scenarios, designed to elicit specific emotions (happiness, anger, sadness, frustration and neutral state).
Annotations were expanded against the original set of classes, to include the identification of nine distinct emotions (angry, excited, fear, sad, surprised, frustrated, happy, disappointed, and neutral), along with metrics for valence, arousal, and dominance.
V-B Experimental setup
As the structure of the database is different from CLSE, we have taken slightly modified approach to verify our hypothesis. The recording sessions were either improvised or scripted, with 12 different improvised scenarios and 3 scripts. We argue that the type of scenario in the recording session represents enough lexical context to the classifier that it can diminish the other factors (prosody etc.).
To verify such claim, we concieved three different experiments. In the first one, we computed the most probable label per given session in a similar fashion to III-B. Then we obtained ASR transcripts with Whisper ASR (large-v3) and incorporated the sentence embeddings using the same extractor as in CLSE experiment. We then applied multi-class logistic regression for topic and emotion label classification. Lastly, we repeated the same with ground truth transcriptions provided by the authors of the database.
All three experiments were conducted in the 5-fold cross-validation setup on a subset of the dataset comprising 5,502 sentences labeled as ’angry’, ’happy’, ’sad’, and ’neutral’. The limited set of emotions is obtained as a standard pre-processing of IEMOCAP database.
VI IEMOCAP:Results
The results of our experiments and also results from comparable speech-based experiments taken from [15] are in Table IV.
TABLE IV: Performance of different systems on IEMOCAP
System UAR
Scenario based 0.62
ASR transcriptions based 0.61
Ground truth transcriptions based 0.61
w2v2 based [16] 0.67
HuBERT based [17] 0.68
MFCCs, Spectrogram, w2v2 based [18] 0.71
These results not only highlight the robustness of textual features in emotion modeling but also invite further inquiry into the diminished role of paralinguistic elements. Intriguingly, our text based experiments yielded only $\sim 9\%$ absolutely worse results than the state-of-the-art large pre-trained models. This emphasizes the need for additional research to explore the complex interplay between textual and paralinguistic features in emotion classification.
VII Discussion and conclusions
This study reveals a critical, often-overlooked aspect of paralinguistic speech recognition: the significant lexical overlap in commonly used datasets. Our analysis of the CLSE and IEMOCAP datasets demonstrates that machine learning models may inadvertently learn text-dependent features rather than the targeted paralinguistic traits. This urges the community to reassess the integrity of current datasets and methodologies.
Reliance on ASR-focused pre-trained models like HuBERT risks conflating lexical and paralinguistic features. While proficient in text-dependent tasks, these models may obfuscate evaluations aimed at paralinguistic recognition, thereby exacerbating lexical overlap issues.
While our results suggest that the lexical features in the speech data significantly influence the classification metrics, they do not negate the importance of paralinguistic features altogether. In the tables, it is clear that when speech features are employed, indeed the performance of systems is better. They merely call for a more careful approach to evaluating paralinguistic recognition systems, with an explicit focus on decoupling textual and non-textual features.
We suggest that future work should focus on the development of methods for reducing text-dependency in existing datasets. Additional evaluations should also be conducted using datasets that have been explicitly designed to minimize lexical overlap, thereby offering a more reliable platform for paralinguistic studies.