### v2

# Beyond the Labels: Unveiling Text-Dependency in Paralinguistic Speech Recognition Datasets

Abstract
Paralinguistic traits such as cognitive load and emotion are recognized as important areas in speech recognition research, often examined through specialized datasets such as CLSE and IEMOCAP. However, the integrity of these datasets is seldom scrutinized for text dependency. This paper critically evaluates the assumption that machine learning models trained on such datasets genuinely learn to identify paralinguistic traits, rather than merely capturing lexical features. By examining the lexical overlap in these datasets and testing the performance of machine learning models, we expose significant text dependency in trait labeling. Our results suggest that some machine learning models, especially large pre-trained models such as HuBERT, might inadvertently focus on lexical characteristics rather than the intended paralinguistic features.
Index Terms:
Paralinguistic Traits, Speech Recognition, Cognitive Load, Emotion Recognition, Lexical Overlap, Machine Learning, Datasets, Text-Dependency
I Introduction
While the primary focus of speech recognition research gravitates towards Automatic Speech Recognition (ASR), the study of paralinguistic traits, such as cognitive load, physiological stress, and emotions, remains a significant field too. These traits are of interest for applications ranging from human-computer interaction to psychological research and rely heavily on dedicated datasets. Two key datasets are Cognitive Load with Speech and EGG (CLSE[1], commonly used for cognitive load recognition, and Interactive Emotional Dyadic Motion Capture (IEMOCAP[2], primarily employed for emotion recognition.
A prevailing assumption is that machine learning algorithms trained on these datasets learn to recognize paralinguistic traits based on observable physiological or psychological changes in speech production. However, this paper challenges this assumption by revealing a critical oversight: we provide evidence of significant lexical correlation between the labels (e.g., cognitive load or emotion) and the uttered sentences within these datasets.
We analyze CLSE and review the design of IEMOCAP to substantiate this claim. Given the implications of our findings, the paper serves as a call reevaluation of existing datasets and methodologies to ascertain that machine learning systems are learning to recognize what they are designed to recognize.
II Related Work
II-A Datasets in Focus
Two key datasets facilitate paralinguistic studies:
  • CLSE (Cognitive Load with Speech and EGG): This is intended to test recognition of cognitive load by incorporating both speech recordings and Electroglottograph (EGG) signals.
  • IEMOCAP (Interactive Emotional Dyadic Motion Capture) [^1] https://sail.usc.edu/iemocap/index.html: It covers about 12 hours of scripted and spontaneous dialogues. It captures speech and facial and hand movements to study a range of emotions.
II-B Machine Learning Approaches to Paralinguistics
Various machine learning algorithms, from classic UBM-iVector [3], through LSTM [4] to recent large pre-trained models like wav2vec [5], have been employed on these datasets. The general presumption is that these models are capturing paralinguistic features rooted in physiological or psychological changes [6, 7].
Pepino et al [8] discovered the text-dependency issue in their work on Emotion Detection on IEMOCAP. In their experiments they observed over-optimistic results while using the original train-test splits. Yuanchao et al [9] analyzed wav2vec performance on IEMOCAP and found that this model prioritizes linguistic content over para-linguistic information.
In paralinguistic research, it is essential to consider both the content and the manner of speech. Physiological and prosodic features often capture the manner of speech, while lexical choices, semantics, and syntax provide insights into the content. Ignoring either aspect may lead to an incomplete understanding of paralinguistic traits. This study emphasizes the importance of acknowledging lexical correlates when analyzing paralinguistic datasets to ensure comprehensive evaluation and accurate model training.
III Hypothesis and Methodology
III-A Hypothesis
Our hypothesis posits that machine learning models trained on paralinguistic speech recognition datasets like CLSE and IEMOCAP may exhibit significant text dependency. This dependency could undermine the models’ ability to genuinely learn paralinguistic traits, instead focusing on lexical characteristics of the speech data.
III-B Methodology
To test this hypothesis, we conducted experiments on both the CLSE and IEMOCAP datasets. We employed various machine learning models, including classic and state-of-the-art approaches, and analyzed their performance with and without lexical overlap. Our methodology involves steps seen in Figure 1.
AudioPreprocessing Whisper ASRSentence Embeddings LaBSEClustering AHCClassification Cluster IdentityEvaluation UARClassification MLREvaluation UARCLSEIEMOCAP
Figure 1: Block scheme of the experiment structure for different datasets
IV Dataset and Task Description
IV-A CLSE
Our primary study focuses on the main part of the CLSE database, termed as CLSE-Span, which employs the  Reading Span task (RSPAN[10]. In RSPAN, participants are asked to validate sentences for their logical coherence while memorizing a sequence of letters. Sentences are randomly chosen from a closed set. The task is divided into sets, where each set contains varying trials of sentence validation and letter memorization. The database consists of 21 such sets with each participant producing 75 utterances, totaling 1800 utterances. The average utterance duration is four seconds. Labels for the cognitive load values are: low, medium, and high load.
IV-B IEMOCAP
The IEMOCAP dataset comprises 151 dialog videos, featuring a pair of speakers in each session, resulting in a total of 302 individual videos with approximately 12 hours of data. Ten actors were recorded in dyadic sessions (five sessions with two subjects each). They were asked to perform three selected scripts with clear emotional content. In addition to the scripts, the subjects were also asked to improvise dialogs in hypothetical scenarios, designed to elicit specific emotions (happiness, anger, sadness, frustration, and neutral state). The original annotations were expanded by the authors after the collection. The new annotated classes contain nine distinct emotions (angry, excited, fear, sad, surprised, frustrated, happy, disappointed, and neutral), along with metrics for valence, arousal, and dominance.
V Experimental Setup
V-1 Data Preprocessing
We used Whisper ASR [11], a pre-trained multilingual large-v3 model, to transcribe the speech data from both datasets. Subsequently, we utilized Sentence Transformer embeddings as features, using the Language-Agnostic BERT Sentence Embeddings (LaBSE) model [12]. LaBSE utilizes a dual-encoder architecture trained in two steps: first on 17 billion monolingual sentences via Masked Language Modeling, and then further refined on six billion translation pairs covering 109 languages. The final model is publicly accessible [^2] https://huggingface.co/sentence-transformers/LaBSE and was used to extract 768-dimensional sentence embeddings.
V-2 Clustering
For the CLSE dataset, we applied Agglomerative Hierarchical Clustering (AHC) to the sentence embeddings to group similar sentences. After manually correcting minor clustering errors, we identified 81 different sentences. Each sentence was assigned an ID based on its cluster membership.
V-3 Classification
Using the clustered data, we assigned the most frequent cognitive load label in the training set to each cluster in the CLSE dataset. These labels were used as a proxy to predict the cognitive load of utterances in the test and validation sets based on their cluster identity, effectively bypassing the use of any machine learning model for classification.
For the IEMOCAP dataset, we analyzed the recording sessions based on the type of scenario (improvised or scripted). We computed the most probable label per session, obtained ASR transcripts with Whisper ASR, and incorporated the sentence embeddings using the LaBSE model. We then applied multi-class logistic regression for topic and emotion label classification, and repeated the experiments using ground truth transcriptions provided by the dataset authors.
V-4 Evaluation
We evaluated the impact of lexical overlap by comparing performance metrics using the Unweighted Average Recall (UAR). This metric offers a balanced assessment of performance across all classes, making it suitable for evaluating imbalanced datasets. We compared the performance of models trained on the original dataset splits with those trained on fixed (shuffled) splits, where lexical redundancy was minimized.
V-A CLSE
We preprocessed, clustered and classified the CLSE dataset as described in previous section.
To further corroborate our hypothesis, we reshuffled the dataset splits based on cluster IDs. This process introduced speaker overlap but did not affect the cognitive load estimation, as each speaker experiences a full range of cognitive loads in the dataset.
V-B IEMOCAP
As the structure of the IEMOCAP database is different from CLSE, we have adopted a slightly modified approach to verify our hypothesis. IEMOCAP sessions are either improvised or scripted, with 12 different improvised scenarios and three scripts. We argue that the scenario of the recording session represents enough lexical context to the classifier that it can diminish the other factors (prosody etc.).
To verify this claim, we conceived three different experiments: In the first one, we computed the most probable label per given session in a similar fashion to Section V-A. Then we obtained ASR transcripts with Whisper ASR (large-v3) and incorporated the sentence embeddings using the same extractor as in the CLSE experiment. We then applied multi-class logistic regression for topic and emotion label classification. Lastly, we repeated the same with ground truth transcriptions provided with the database.
All three experiments were conducted in five-fold cross-validation setup on a subset of the dataset comprising 5,502 sentences labeled as ’angry’, ’happy’, ’sad’, and ’neutral’. The limited set of emotions is obtained as a standard pre-processing of IEMOCAP and corresponds to the author’s recommendations [13].
VI Results
VI-A CLSE
We assessed the influence of lexical overlap on classification performance using both original and fixed (shuffled) CLSE data splits. The original CLSE data displayed good performance metrics, but revealed patterns in the validation set that confirmed the lexical factors’ role.
TABLE I: Performance in Original vs Fixed CLSE Splits for Cluster-Based Classification
Original CLSE Fixed CLSE
Train 0.67 0.33
Validation 0.79 0.33
Test 0.57 0.33
After minimizing lexical redundancy in shuffled splits, the performance metrics notably decreased to the chance level, as can be seen in Table I, further substantiating the lexical overlap’s impact on classification.
We replicated the UBM-iVector system from the ComPaRE 2014 [14] challenge, using a 64-component Universal Background Model (UBM) with a 50-dimensional iVector extractor and Support Vector Machine (SVM) for classification [3]. Using the same approach as in the cluster-based approach, we obtained results showing significant degradation of the performance on validation and test sets (Table II).
TABLE II: Performance in Original vs Fixed CLSE Splits for ComPaRE2014 System
Split Original CLSE Fixed CLSE
Train 0.94 0.98
Validation 0.75 0.51
Test 0.64 0.51
To corroborate our findings with a more recent speech-based system, we used HuBERT [15] (version large-ll60k) pre-trained on 60k hours from the Libri-Light dataset [16] as a feature extractor. As a classifier, we used an Attentive correlation pooling layer from [17]. This classifier uses a multi-head attention layer, where attention weights resemble a mixture model with heads parametrizing the mixture component. The results of our experiments are shown in Table III. They again manifest the same type of behavior as the previous experiments.
TABLE III: Performance in Original vs Fixed CLSE Splits for HuBERT Based System
Split Original CLSE Fixed CLSE
Train 0.82 0.76
Validation 0.76 0.64
Test 0.74 0.52
VI-B IEMOCAP
For IEMOCAP, the results of our experiments and comparable speech-based experiments taken from [18] are shown in Table IV.
TABLE IV: Performance of Different Systems on IEMOCAP
System UAR
Scenario based 0.62
ASR transcriptions based 0.61
Ground truth transcriptions based 0.61
w2v2 based [19] 0.67
HuBERT based [20] 0.68
MFCCs, Spectrogram, w2v2 based [21] 0.71
[…]
VII Discussion and Conclusions
This study reveals a critical, often-overlooked aspect of paralinguistic speech recognition: the significant lexical overlap in commonly used datasets. Our analysis of the CLSE and IEMOCAP datasets demonstrates that machine learning models may inadvertently learn text-dependent features rather than the targeted paralinguistic traits. This urges the community to reassess the integrity of current datasets and methodologies.
Reliance on ASR-focused pre-trained models like HuBERT risks conflating lexical and paralinguistic features. While excelling in text-dependent tasks, these models may obfuscate evaluations aimed at paralinguistic recognition.
While our results suggest that the lexical features in the speech data significantly influence the classification metrics, they do not negate the importance of paralinguistic features altogether. It is clear that when speech features are employed, indeed the performance of systems is better. They merely call for a more careful approach to evaluating paralinguistic recognition systems, with an explicit focus on decoupling textual and non-textual features.
We suggest that future work should focus on the development of methods for reducing text dependency in existing datasets. Additional evaluations should also be conducted using datasets that have been explicitly designed to minimize lexical overlap.